nolegirl24:

A firecracker shoots up from a hill 160 feet high, with an initial speed of 90 feet per second. Using the formula H(t) = −16t^2 + vt + s, approximately how long will it take the firecracker to hit the ground? Five seconds Six seconds Seven seconds Eight seconds

dude:

We need to substitute values that are given to us by the context v is velocity, what is the speed given to us here?

nolegirl24:

90 ft per second

dude:

\(H(t) = −16t^2 + vt + s\) we substitute v for 90 \(H(t) = −16t^2 + 90t + s\) The s is the original height, what was the original height given to us here?

nolegirl24:

Original height would be 160 ft high

dude:

Yes So our equation is: \(H(t) = −16t^2 + 90t + 160\) When it says when the firecracker reaches the ground, we want the height to be 0 (because its in the ground) We set this equal to 0 \(0 = −16t^2 + 90t + 160\)

nolegirl24:

okay got it down

dude:

Do you know how to solve for t?

nolegirl24:

I have no idea what t would end up being

dude:

Okay we can use the quadratic equation, do you know what that is?

nolegirl24:

I am supposed to but i honestly forgot and i cant find my notebook

nolegirl24:

I am using a spare notebook until i can find the one from yesterday

dude:

\(t=\large \frac{-b\pm\sqrt{b^2-4ac}}{2a}\) Have you seen this before?

nolegirl24:

o.o ive seen it but never figured out how to use it

dude:

Okay so we need to find the a, b and c values for them A quadratic equation is written as \(y=ax^2+bx+c\) With what we have \(H(t)=-16t^2+90t+160\) What are the a, b and c

nolegirl24:

A: -16 B: 90 C:160 Right.....?

dude:

Yes

dude:

Now we substitute what we have into this \(a=-16\) \(b=90\) \(c=160\) \(t=\large \frac{-b\pm\sqrt{b^2-4ac}}{2a}\)

nolegirl24:

Alright

nolegirl24:

So now we just substitute them?

dude:

Yes

nolegirl24:

Okay one sec

nolegirl24:

Idk how to do the thing that you did with the equation so i can show you.....

dude:

`\(t=\large \frac{- b \pm\sqrt{b ^2-4 a c}}{2 a}\)` Copy that and insert the a,b and c

nolegirl24:

\(t=\large \frac{- -90 \pm\sqrt{90 ^2-4 (-16) (160)}}{2 -16}\)

dude:

Yes, double negatives will make a positive --90 = 90 And multiply what is in within the root Multiply the 2 and -16

nolegirl24:

-32

dude:

Right so we have \(t=\large \frac{90 \pm\sqrt{90 ^2-4 (-16) (160)}}{-32}\) Multiply what is inside the square root now \(t=\large \frac{90 \pm\sqrt{\color{red}{90 ^2-4 (-16) (160)}}}{-32}\)

nolegirl24:

okay one sec

nolegirl24:

so like 90^2 = 8100? and would i do -4 * -16 and -4 * 160?

dude:

Yes \(90^2\) is 8100 and then you multiply -4 times -16 and then the answer between those times 160

nolegirl24:

So 64-160?

dude:

\(64 \times 160\)

nolegirl24:

Ohh okay one sec

nolegirl24:

so that would equal 10240

dude:

Right Now add 8100 and 10240

nolegirl24:

18340

dude:

Right so we have \(\large \frac{90\pm\sqrt{18340}}{-32}\) Find \(\sqrt{18340}\) using your calculator

nolegirl24:

thats square root of 18340?

dude:

Yes

nolegirl24:

135.425256138

dude:

You can just round it do \(135.425\) Basically now your answer can be: \(\large \frac{90+135.425}{-32}\) or \(\large \frac{90-135.425}{-32}\) Solve both

nolegirl24:

225.425/-32 or -45.425/-32

dude:

Yes, divide them

nolegirl24:

okay one sec

nolegirl24:

-7.044 or 1.419

dude:

Ah made a sign mistake \(\large \frac{-90+135.425}{-32}\) or \(\large \frac{-90-135.425}{-32}\) Those answers are right, the signs are just flipped \(7.04~and~-1.419\)

dude:

Okay so from here the question asks for time We cannot have negative time so \(-1.419\) is deleted \(7.04\) is close to 7 seconds so 7 is your answer

nolegirl24:

Ohh okay thank you. You really have been helping me alot.

dude:

Of course, thats what I am here for

nolegirl24:

^~^

dude:

:)